# DermaCon-IN: A Multi-concept Annotated Dermatological Image Dataset of Indian Skin Disorders for Clinical AI Research

### v1

@@ Abstract
Artificial intelligence is poised to augment dermatological care by enabling scalable image-based diagnostics. Yet, the development of robust and equitable models remains hindered by datasets that fail to capture the clinical and demographic complexity of real-world practice. This complexity stems from region-specific disease distributions, wide variation in skin tones, and the underrepresentation of outpatient scenarios from non-Western populations. We introduce DermaCon-IN, a prospectively curated dermatology dataset comprising over 5,450 clinical images from approximately 3,000 patients across outpatient clinics in South India. Each image is annotated by board-certified dermatologists with over 240 distinct diagnoses, structured under a hierarchical, etiology-based taxonomy adapted from Rook’s classification. The dataset captures a wide spectrum of dermatologic conditions and tonal variation commonly seen in Indian outpatient care. We benchmark a range of architectures—including convolutional models (ResNet, DenseNet, EfficientNet), transformer-based models (ViT, MaxViT, Swin), and Concept Bottleneck Models to establish baseline performance and explore how anatomical and concept-level cues may be integrated. These results are intended to guide future efforts toward interpretable and clinically realistic models. DermaCon-IN provides a scalable and representative foundation for advancing dermatology AI in real-world settings.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a promising solution for enhancing dermatological diagnosis and triage, particularly in resource-constrained regions with limited access to dermatologists. However, a critical bottleneck remains: the lack of representative training data that adequately captures diversity in disease presentations and skin tones based on regional relevance [4]. Most AI models for dermatology to date have been developed and benchmarked using datasets predominantly sourced from North American, European, or Australasian populations [1, 25]. This geographic skew has introduced performance biases, especially for underrepresented groups such as individuals, patients presenting with diseases more prevalent outside Western contexts, and with darker skin tones.
@@
To address these limitations, we introduce a new dermatology image dataset curated from Indian outpatient clinics. To the best of our knowledge, it is the first densely annotated dataset centered around Indian skin phototype and is designed to improve diversity in both disease coverage and skin tone representation for dermatological AI research. It complements existing datasets by capturing the phenotypic and pathological landscape of a population historically underserved in global medical AI efforts. In addition, the dataset also aims to support explainable modelling by reflecting how dermatologists diagnose, through the combined use of anatomical location and visual descriptors. The Key contributions of this work are as follows:
@@ 2 Related Work
Table 1: Comparative Survey of existing Dermatology Datasets available for AI research [Columns: A: Neoplasm & Tumors Centric, B: Broad Skin Disease Spectrum, C: Dermoscopic Single Lesion Focus, D: Real-time Multi-lesion Multi-Focus, E: Body Part, F: Lesion Descriptor]
@@
A B C D E F Web Scraped (Atlas) Geographic Location Present #Images Hierarchical #level[#count]
@@
In South Asia, outpatient dermatology is dominated by inflammatory, infectious, pigmentary, and appendageal diseases [39], yet remains underrepresented in existing datasets. We address this gap with a dataset of $5{,}450$ high-resolution clinical images collected prospectively from $\sim$3,000 Indian patients. It covers the disease spectrum aligned with Indian and global burden data. Each image retains anatomical context and is annotated by board-certified dermatologists with standardised diagnosis, Fitzpatrick and MST skin tone ratings, distribution of which is shown in Figure  2. Unlike prior work such as SkinCon [5], which retrofitted a set of lesion descriptors onto existing datasets, we capture both lesion and anatomical concepts at source and leverage the full concept set for statistical validation and model benchmarking.
@@ 3.2 Image Acquisition Protocol
Image capture was designed to mirror real-world dermatology workflows, emphasizing both lesion detail and the broader anatomical region. Photographs were taken using high-resolution smartphone devices—108MP Android, 48MP iPhone Pro, and 12–36MP cameras—under ambient clinical lighting. Images include the affected body part with surrounding skin to preserve anatomical context and support spatial modeling. A standardized protocol guided acquisition across sites, allowing relevant variations in angle, distance, and lighting to reflect clinical realism, unlike prior datasets focused on dermoscopic or tightly cropped views.
@@ 3.4 Annotation Process
Figure 3: Cohen’s Kappa scores for cross validation of annotations provided by $4$ doctors (D1, D2, D3, D4). DA_DB is a cumulative score.
@@ 3.6 Final Dataset Composition
The final composition contains $5{,}450$ high-resolution JPEG images across $8$ top-level etiologic classes, $19$ clinically meaningful subclasses, and $245$ fine-grained disease labels. Each sample includes dense metadata: hierarchical disease labels, body parts, skin lesion descriptors, Fitzpatrick [7] and Monk skin tone [31] scores, diagnostic certainty, and image gradability. We consider $49$ body parts and $47$ lesion descriptors as concepts, which account for $96$ unique concepts.
@@ 4 Dataset Overview and Statistics
DermaCon-IN is organized around an etiologically informed, clinically validated taxonomy rooted in Rook’s [9] classification and aligned with ICD-11 [41]. The dataset includes $8$ high-level diagnostic categories ranging from Infectious to Neoplastic, including No Definite Diagnosis — reflecting the full spectrum of dermatological conditions prevalent in South Asian outpatient settings [39]. These main classes are further expanded into 19 subclasses to capture hybrid and co-occurring disease states. For example, subclass combinations such as Inflammatory + Infectious (Bacterial) reflect polymorphic real-world presentations of superimposed two diseases. Each top-level category is populated with sufficient training instances for stratified benchmarking; notably, high-burden groups like fungal infections and eczema dominate in volume, while rare but clinically significant categories—e.g., keratinization disorders remain represented with enough density to enable few-shot generalization. The Disease label reflects $245$ distinct disease types, which follow a long-tailed distribution with a log-normal fit to leaf-node frequencies, yielding an exponent of $1.8$, as expected in real-world clinical settings. To better understand label diversity in the dataset, refer Figure 4. This visualization underscores the need for imbalance-aware learning strategies.
@@
The visual descriptors of skin lesions annotated in DermaCon-IN follow an established dermatological vocabulary commonly used by clinicians to describe skin lesions in line with extant clinical literature [3]. Inspired by frameworks [5] previously validated in dermatology AI literature, we consulted board-certified dermatologists to select clinically meaningful descriptors aligned with disease patterns frequently encountered in Indian outpatient settings. Rather than assigning fixed concepts per disease category, each image was independently annotated, reflecting the realistic variability in lesion appearance across patients, body regions, and skin types. Statistical analysis in Figure 5 of concept-disease correlations validated this flexible annotation strategy that supports concept-supervised learning approaches. Details of Descriptors utilised for annotation are provided in Datasheet (Supplementary Material).
Figure 5: Chord diagrams depicting the most informative concept-class correlations for each main skin disease class. (A) Correlations between disease classes and clinical descriptors (e.g., morphology, texture, pigmentation); (B) correlations between disease classes and annotated body regions. For both, Pearson correlation coefficients were computed, and the top-$k$ positive and bottom-$k$ negative associations per class were selected. Ribbon widths represent correlation magnitude, and color encodes direction and strength (blue: positive, red: negative), highlighting characteristic clinical and anatomical patterns across disease groups.
@@
Such distributional cues are vital for accurate real-world diagnosis but are often lost in lesion-only images. Our dataset includes annotations for $49$ distinct body sites, organized using both precise clinical terms and broader regional tags (e.g., trunk, extremities, scalp). A List of body parts annotated is provided in Datasheet (Supplementary Material).
@@
We examined Pearson correlation coefficients between (a) Disease descriptors and disease categories, and (b) anatomical body regions and disease categories, (refer Figure 5) to assess the biological plausibility and interpretability of our clinical annotations. These chord diagrams reveal statistically meaningful associations that align well with established dermatological knowledge. For instance, positive (high) correlations are observed between erythema, vesicle, and scale with inflammatory disorders, and between hyperkeratotic plaques and keratinisation disorders. Similarly, white patches and pigmented lesions show high positive associations with pigmentary and infectious disorders, respectively, underscoring the dermatologic specificity of the disease descriptors in general. Anatomical correlations further reinforce clinical fidelity. Keratinisation disorders predominantly localize to the soles and palms, consistent with plantar keratoderma patterns. Skin appendageal disorders, such as acne and seborrheic dermatitis, strongly associate with sebaceous-rich zones like the scalp and cheeks [3, 30]. These findings validate the anatomical tropisms encoded in the dataset.
@@ 5 Challenges, Opportunities & Limitations
The dataset presents a range of challenges and opportunities that stem from the inherent complexity of real-world clinical data, offering both practical constraints and avenues for robust model development:
@@
Images were cropped post-acquisition to remove garments and background clutter where feasible, though incidental artifacts (e.g., jewelry) remain in some cases. The resulting variability in resolution, arising from diverse capture devices and aspect ratios, is a natural outcome, but advantageous. It reflects real-world conditions, where patients may crop or capture images themselves, and encourages model robustness to such variations. Image heights range from $296$ to $4{,}608$ pixels and widths from $346$ to $4{,}608$ pixels, with a mean resolution of $2{,}300$x2,05720572{,}0572 , 057 pixels. The average image area is $4.97$M pixels ($\pm$3.01M), with an interquartile range of $2.62$M–$6.69$M pixels (Figure 6), indicating consistently high-fidelity input for fine-grained modeling.
@@
  • Long-tailed Disease Labels: The dataset embraces the inherent long-tailed nature of dermatological diagnoses, where few diseases are prevalent and many are rare. This structure presents a valuable opportunity to train models that generalize across the full clinical spectrum, including rare disease categories.
@@
  • Long-tailed distribution: The natural skew in concept frequencies mirrors real-world prevalence, where rare but critical findings coexist with common patterns. This creates opportunities to tackle challenges in multi-concept learning and rare concept detection—core problems in clinical AI.
@@
In accordance with ethical considerations, all images were anonymised by masking identifiable facial features, such as the eyes, and cropping facial regions where necessary. While essential for protecting patient identity, this may limit the model’s ability to accurately learn or detect diseases that primarily manifest on the face.
@@ 6 Benchmarking with Models
Table 2: Performance benchmarking of the proposed dataset on the $8$-main class diagnosis classification task, conducted using both CNN- and ViT-based standard architectures. All metrics are averaged over $5$ random seeds.
Model Pre-trained Accuracy Precision Sensitivity F1 Score
ResNet50 [13] - $47.45_{\pm 0.40}$ $46.59_{\pm 0.51}$ $47.44_{\pm 0.41}$ $46.43_{\pm 0.60}$
DenseNet121 [16] - $49.30_{\pm 0.30}$ $47.49_{\pm 0.86}$ $49.30_{\pm 0.29}$ $48.17_{\pm 0.74}$
ResNet50 [13] ImageNet $64.31_{\pm 0.22}$ $63.41_{\pm 0.29}$ $64.31_{\pm 0.23}$ $63.31_{\pm 0.25}$
DenseNet121 [16] ImageNet $65.20_{\pm 0.48}$ $64.62_{\pm 0.45}$ $65.20_{\pm 0.48}$ $64.37_{\pm 0.35}$
EffNet-B4_b4 [35] ImageNet $64.28_{\pm 0.34}$ $63.53_{\pm 0.64}$ $64.27_{\pm 0.34}$ $63.38_{\pm 0.39}$
ViT-B/16-224 [6] ImageNet $64.09_{\pm 1.03}$ $62.59_{\pm 1.03}$ $62.88_{\pm 1.67}$ $62.98_{\pm 1.02}$
ViT-B/16-384 [6] ImageNet $66.95_{\pm 0.19}$ $65.39_{\pm 0.13}$ $66.95_{\pm 0.20}$ $65.78_{\pm 0.06}$
MaxViT-B/512 [38] ImageNet $66.92_{\pm 0.48}$ $66.30_{\pm 0.80}$ $66.92_{\pm 0.48}$ $65.90_{\pm 0.73}$
Swin-B/4W12-384 [24] ImageNet $70.41_{\pm 0.41}$ $69.83_{\pm 0.37}$ $70.41_{\pm 0.42}$ $69.69_{\pm 0.46}$
Table 3: Performance comparison of model variants using the Swin-B/4W12-384 backbone. The first two rows are baselines without a concept bottleneck (CB) layer, used for $8$-main class (MC) and $19$-subclass (SC) classification. The next four rows report CBMs trained with different concept sets: lesion descriptors ($47$), body parts ($49$), and both ($96$). The last merged rows present individual layer (SC & MC) performance of a Hierarchical CBM (Type 1 & 2) combining the CB layer with joint SC and MC classification as described in Fig  7. All metrics are results of end-to-end training and are averaged over $5$ random seeds.
@@
Concepts + SC Descriptors & Body parts $55.54_{\pm 0.73}$ $56.85_{\pm 0.92}$ $55.53_{\pm 0.76}$ $55.75_{\pm 0.85}$ $79.42_{\pm 0.17}$
(Type1) Concepts + SC Descriptors & Body parts $53.98_{\pm 0.40}$ $56.12_{\pm 0.74}$ $53.98_{\pm 0.39}$ $54.49_{\pm 0.66}$ $76.13_{\pm 1.52}$
+ MC $67.78_{\pm 0.47}$ $67.32_{\pm 0.67}$ $67.66_{\pm 0.48}$ $66.92_{\pm 0.37}$ $79.53_{\pm 0.62}$
[…]
+ MC $69.90_{\pm 0.20}$ $68.82_{\pm 0.36}$ $69.89_{\pm 0.19}$ $69.08_{\pm 0.31}$ $77.01_{\pm 2.24}$

### v2

@@ Abstract
Artificial intelligence is poised to augment dermatological care by enabling scalable image-based diagnostics. Yet, the development of robust and equitable models remains hindered by datasets that fail to capture the clinical and demographic complexity of real-world practice. This complexity stems from region-specific disease distributions, wide variation in skin tones, and the underrepresentation of outpatient scenarios from non-Western populations. We introduce DermaCon-IN, a prospectively curated dermatology dataset comprising 5,450 clinical images from 3,002 patients across outpatient clinics in South India. Each image is annotated by board-certified dermatologists with 245 distinct diagnoses, structured under a hierarchical, aetiology-based taxonomy adapted from Rook’s classification. The dataset captures a wide spectrum of dermatologic conditions and tonal variation commonly seen in Indian outpatient care. We benchmark a range of architectures, including convolutional models (ResNet, DenseNet, EfficientNet), transformer-based models (ViT, MaxViT, Swin), and Concept Bottleneck Models to establish baseline performance and explore how anatomical and concept-level cues may be integrated. These results are intended to guide future efforts toward interpretable and clinically realistic models. DermaCon-IN provides a scalable and representative foundation for advancing dermatology AI.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a promising solution for enhancing dermatological diagnosis and triage, particularly in resource-constrained regions with limited access to dermatologists. However, a critical bottleneck remains: the lack of representative training data that adequately captures diversity in disease presentations and skin tones based on regional relevance [4]. Most AI models for dermatology to date have been developed and benchmarked using datasets predominantly sourced from North American, European, or Australasian populations [1, 26]. This geographic skew has introduced performance biases, especially for underrepresented groups such as individuals, patients presenting with diseases more prevalent outside Western contexts, and those with darker skin tones.
@@
To address these limitations, we introduce a new dermatology image dataset curated from Indian outpatient clinics. To the best of our knowledge, it is the first densely annotated dataset centered around Indian skin phototype and is designed to improve diversity in both disease coverage and skin tone representation for dermatological AI research. It complements existing datasets by capturing the phenotypic and pathological landscape of a population historically underserved in global medical AI efforts. In addition, the dataset also aims to support explainable modeling by reflecting how dermatologists diagnose, through the combined use of anatomical location and visual descriptors. The Key contributions of this work are as follows:
@@ 2 Related Work
Table 1: Comparative Survey of existing Dermatology Datasets available for AI research [Columns: A: Neoplasm & Tumors Centric, B: Broad Skin Disease Spectrum, C: Dermoscopic Single Lesion Focus, D: Real-time Multi-lesion Multi-Focus, E: Body Part, F: Lesion Descriptor, G: Rook’s classification labels]
@@
A B C D E F G Web Scraped (Atlas) Geographic Location Present #Images Hierarchical #level[#count]
@@
In South Asia, outpatient dermatology is dominated by inflammatory, infectious, pigmentary, and appendageal diseases [42], yet remains underrepresented in existing datasets. We address this gap with a dataset of 5,4505{,}450 high-resolution clinical images collected prospectively from 3,002 Indian patients. It covers the disease spectrum aligned with Indian and global burden data. Each image retains anatomical context and is annotated by board-certified dermatologists with standardised diagnosis, as well as Fitzpatrick and MST skin tone ratings, which align with patterns observed in the Indian context [32, 33, 15]. The distribution of which is shown in Figure  2. Unlike prior work such as SkinCon [5], which retrofitted a set of lesion descriptors onto existing datasets, we capture both lesion and anatomical concepts at source and leverage the full concept set for statistical validation and model benchmarking.
@@ 3.2 Image Acquisition Protocol
Image capture was designed to mirror real-world dermatology workflows, emphasizing both lesion detail and the broader anatomical region. Photographs were taken using high-resolution smartphone devices, viz, 108MP Android, 48MP iPhone Pro, and 12–36MP cameras, under ambient clinical lighting. Images include the affected body part with surrounding skin to preserve anatomical context and support spatial modeling. A standardized protocol guided acquisition across sites, allowing relevant variations in angle, distance, and lighting to reflect clinical realism, unlike prior datasets focused on dermoscopic or tightly cropped views.
@@ 3.4 Annotation Process
Figure 3: Cohen’s Kappa scores for cross validation of annotations provided by 44 doctors (D1,D2,D3,D4)
@@ 3.6 Final Dataset Composition
The final composition contains 5,4505{,}450 high-resolution JPEG images across $8$ top-level etiologic classes, $19$ clinically meaningful subclasses, and $245$ fine-grained disease labels. Each sample includes dense metadata: hierarchical disease labels, body parts, skin lesion descriptors, Fitzpatrick [7] and Monk skin tone [34] scores, diagnostic certainty, and image gradability. We consider $49$ body parts and $47$ lesion descriptors as concepts, which account for $96$ unique concepts.
@@ 4 Dataset Overview and Statistics
DermaCon-IN is organized around an aetiologically informed, clinically validated taxonomy rooted in Rook’s [9] classification and aligned with ICD-11 [44]. The dataset includes $8$ high-level diagnostic categories ranging from Infectious to Neoplastic, including No Definite Diagnosis, reflecting the full spectrum of dermatological conditions prevalent in South Asian outpatient settings [42]. These main classes are further expanded into 19 subclasses to capture hybrid and co-occurring disease states. For example, subclass combinations such as Inflammatory + Infectious (Bacterial) reflect polymorphic real-world presentations of superimposed two diseases. Each top-level category is populated with sufficient training instances for stratified benchmarking; notably, high-burden groups like fungal infections and eczema dominate in volume, while rare but clinically significant categories, e.g., keratinisation disorders remain represented with enough density to enable few-shot generalisation. The Disease label reflects $245$ distinct disease types, which follow a long-tailed distribution with a log-normal fit to leaf-node frequencies, yielding an exponent of $1.8$, as expected in real-world clinical settings. To better understand label diversity in the dataset, refer Figure 4. This visualization underscores the need for imbalance-aware learning strategies.
@@
The visual descriptors of skin lesions
Figure 5: Chord diagrams of Pearson correlations ($r$) between eight disease classes and two concept types: (A) lesion descriptors and (B) body regions. Ribbon width represents within-class correlation strength, and colour encodes direction and magnitude (dark blue = positive; red = negative). Numbers on ribbons indicate $r$. Top-$k$ positive and negative associations per class highlight distinctive clinical and anatomical patterns across disease groups.
annotated in DermaCon-IN follow an established dermatological vocabulary commonly used by clinicians to describe skin lesions in line with extant clinical literature [3]. Inspired by frameworks [5] previously validated in dermatology AI literature, we consulted board-certified dermatologists to select clinically meaningful descriptors aligned with disease patterns frequently encountered in Indian outpatient settings. Rather than assigning fixed concepts per disease category, each image was independently annotated, reflecting the realistic variability in lesion appearance across patients, body regions, and skin types. Statistical analysis in Figure 5 of concept-disease correlations validated this flexible annotation strategy that supports concept-supervised learning approaches. Details of Descriptors utilized for annotation are provided in the Datasheet (Supplementary Material).
@@
Such distributional cues are vital for accurate real-world diagnosis but are often lost in lesion-only images. Our dataset includes annotations for $49$ distinct body sites, organised using both precise clinical terms and broader regional tags (e.g., trunk, extremities, scalp). A List of body parts annotated is provided in the the Datasheet (Supplementary Material).
@@
We examined Pearson correlation coefficients between (a) Disease descriptors and disease categories, and (b) anatomical body regions and disease categories, (refer Figure 5) to assess the biological plausibility and interpretability of our clinical annotations. Each chord diagram visualizes these relationships, where ribbon width indicates the strength of association between a concept and the class (within-class correlation) and ribbon color (dark blue $\rightarrow$ red) encodes the strength of correlation across classes (dark blue = strong positive; red = negative). Numeric labels on ribbons denote the actual correlation coefficients.
For instance, under Pigmentary Disorders, the descriptor White shows a moderately wide ribbon and dark-blue color ($r=+0.71$), reflecting a strong and distinctive association both within the class and relative to other classes. In contrast, Hyperkeratotic plaques under Keratinization Disorders display a wider but lighter-blue ribbon ($r=+0.47$), suggesting it is more class-specific but less distinctive across classes. Overall, the chord diagrams reveal statistically meaningful associations that align well with established dermatological knowledge. For instance, positive (high) correlations are observed between erythema, vesicle, and scale with inflammatory disorders, and between hyperkeratotic plaques and keratinisation disorders. Similarly, white patches and pigmented lesions show high positive associations with pigmentary and infectious disorders, respectively, underscoring the dermatologic specificity of the disease descriptors in general. Anatomical correlations further reinforce clinical fidelity. Keratinisation disorders predominantly localize to the soles and palms, consistent with plantar keratoderma patterns. Skin appendageal disorders, such as acne and seborrheic dermatitis, are strongly associated with sebaceous-rich zones like the scalp and cheeks [3, 31]. These findings validate the anatomical tropisms encoded in the dataset.
@@ 5 Challenges, Opportunities & Limitations
The dataset presents a range of challenges and opportunities that stem from the inherent complexity of real-world clinical data, offering practical constraints and avenues for robust model development:
@@
Images were cropped post-acquisition to remove garments and background clutter where feasible, though incidental artifacts (e.g., jewelry) remain in some cases. The resulting variability in resolution, arising from diverse capture devices and aspect ratios, is a natural outcome, but advantageous. It reflects real-world conditions, where patients may crop or capture images themselves, and encourages model robustness to such variations. Image heights range from $296$ to 4,6084{,}608 pixels and widths from $346$ to 4,6084{,}608 pixels, with a mean resolution of 2,3002{,}300x2,0572{,}057 pixels. The average image area is 4.974.97M pixels ($\pm$3.01M), with an interquartile range of 2.622.62M–6.696.69M pixels (Figure 6), indicating consistently high-fidelity input for fine-grained modeling.
@@
  • Long-tailed Disease Labels: The dataset embraces the inherent long-tailed nature of dermatological diagnoses, where few diseases are prevalent and many are rare. This structure presents a valuable opportunity to train models that generalise across the full clinical spectrum, including rare disease categories.
  • Multi-disease co-occurrences: A subset shows concurrent lesions from multiple disease types on the same anatomical site (e.g., Inflammatory+Fungal, Fungal+Bacterial). We represent these as dedicated subclasses to help models disentangle overlapping pathologies for the main-class label we follow dermatologists’ treatment-priority logic, e.g., an infected eczema is assigned to the Infectious main class (not Inflammatory) because initial management targets the infection. Such cases, though less frequent, are observed in clinical settings, and our targeted misclassification analysis for these samples is provided in the Supplementary Sec. B2.
@@
  • Long-tailed distribution: The natural skew in concept frequencies mirrors real-world prevalence, where rare but critical findings coexist with common patterns. This creates opportunities to tackle challenges in multi-concept learning and rare concept detection, core problems in clinical AI.
@@
In accordance with ethical considerations, all images were anonymized by masking identifiable facial features, such as the eyes, and cropping facial regions where necessary. While essential for protecting patient identity, this may limit the model’s ability to accurately learn or detect diseases that primarily manifest on the face.
@@ 6 Benchmarking with Models
Table 2: Performance benchmarking of the proposed dataset on the 88-main class diagnosis classification task, conducted using both CNN- and ViT-based standard architectures. All metrics are averaged over 55 random seeds.
Model Pre-trained Accuracy Balanced Acc. Precision Sensitivity F1 Score
ResNet50 [13] - $47.45_{\pm 0.40}$ $23.93_{\pm 0.01}$ $46.59_{\pm 0.51}$ $47.44_{\pm 0.41}$ $46.43_{\pm 0.60}$
DenseNet121 [17] - $49.30_{\pm 0.30}$ $25.31_{\pm 0.01}$ $47.49_{\pm 0.86}$ $49.30_{\pm 0.29}$ $48.17_{\pm 0.74}$
ResNet50 [13] ImageNet $64.31_{\pm 0.22}$ $38.77_{\pm 1.40}$ $63.41_{\pm 0.29}$ $64.31_{\pm 0.23}$ $63.31_{\pm 0.25}$
DenseNet121 [17] ImageNet $65.20_{\pm 0.48}$ $37.31_{\pm 0.01}$ $64.62_{\pm 0.45}$ $65.20_{\pm 0.48}$ $64.37_{\pm 0.35}$
EffNet-B4 [38] ImageNet $64.28_{\pm 0.34}$ $35.58_{\pm 0.01}$ $63.53_{\pm 0.64}$ $64.27_{\pm 0.34}$ $63.38_{\pm 0.39}$
ViT-B/16-224 [6] ImageNet $64.09_{\pm 1.03}$ $34.56_{\pm 0.01}$ $62.59_{\pm 1.03}$ $62.88_{\pm 1.67}$ $62.98_{\pm 1.02}$
ViT-B/16-384 [6] ImageNet $66.95_{\pm 0.19}$ $35.78_{\pm 0.02}$ $5.39_{\pm 0.13} 66$$6.95_{\pm 0.20}$ $65.78_{\pm 0.06}$
MaxViT-B/512 [41] ImageNet $66.92_{\pm 0.48}$ $36.07_{\pm 0.01}$ $6.30_{\pm 0.80} 66$$6.92_{\pm 0.48}$ $65.90_{\pm 0.73}$
Swin-B/4W12-384 [25] ImageNet $70.41_{\pm 0.41}$ $45.06_{\pm 0.02}$ $69.83_{\pm 0.37}$ $70.41_{\pm 0.42}$ $69.69_{\pm 0.46}$
Table 3: Performance comparison of model variants using the Swin-B/4W12-384 backbone. The first two rows are baselines without a concept bottleneck (CB) layer, used for 88-main class (MC) and 1919-subclass (SC) classification. The next four rows report CBMs trained with different concept sets: lesion descriptors ($47$), body parts ($49$), and both ($96$). The last merged rows present individual layer (SC & MC) performance of a Hierarchical CBM (Type 1 & 2) combining the CB layer with joint SC and MC classification as described in Fig  7. All metrics are results of end-to-end training and are averaged over 55 random seeds.
@@
(CBM-D) Concepts + SC Descriptors $6.42_{\pm 0.01} 55$$5.72_{\pm 0.01}$ $6.51_{\pm 0.01} 55$$5.63_{\pm 0.01}$ $80.16_{\pm 0.01}$
(CBM-B) Concepts + SC Body parts $5.88_{\pm 0.01} 55$$47_{\pm 0.01} 55.$$5.90_{\pm 0.01}$ $54.87_{\pm 0.01}$ $78.94_{\pm 0.01}$
Concepts + SC Descriptors & Body parts $57.37_{\pm 0.01}$ $57.67_{\pm 0.01}$ $57.27_{\pm 0.01}$ $56.90_{\pm 0.01}$ $78.39_{\pm 0.02}$
(Type1) SC Descriptors & Body parts $53.98_{\pm 0.40}$ $56.12_{\pm 0.74}$ $53.98_{\pm 0.39}$ $54.49_{\pm 0.66}$ $76.13_{\pm 1.52}$
(Type1) MC Descriptors & Body parts $67.78_{\pm 0.47}$ $67.32_{\pm 0.67}$ $7.66_{\pm 0.48} 66$$6.92_{\pm 0.37}$ $79.53_{\pm 0.62}$
(Type 2) SC Descriptors & Body parts $6.11_{\pm 0.67} 55$$5.49_{\pm 0.0.62}$ $56.09_{\pm 0.9}$ $54.49_{\pm 0.66}$ $76.13_{\pm 1.52}$
(Type 2) MC Descriptors & Body parts $69.90_{\pm 0.20}$ $68.82_{\pm 0.36}$ $69.89_{\pm 0.19}$ $69.08_{\pm 0.31}$ $77.01_{\pm 2.24}$